From the given information


Amount in Lakhs


cost of goods sold 580


opening stock 40


closing stock 70


creditors at the beginning of the year 60


creditors at the end of the year 100


cash purchases 45


Original cost of equipment sold 400


Gain on the equipment sold 50


Accumulated depreciation on the equipment 80


Calculate:


a. Total purchases, credit purchases and payment to creditors (5 Marks)


b. Define the term Net book value, Accumulated depreciation calculate the net book value


and cash proceeds from sale of investment (5 Marks)

1
Expert's answer
2022-03-28T14:44:15-0400

a)

Total purchases= closing stock- Opening stock +cost of goods sold

70-40+580= 610

Credit purchases= creditor closing balance+ cash paid- opening creditor balance

= 100+45- 60= 85

Payment to creditor

= Cost of goods sold+ increase in inventor- increase in accounts payable

580+ 30-40= 570

b)

Net book value- is the amount at which an organization records an asset in its accounting records.

Accumulated depreciation is the cumulative depreciation of an asset up to a single point in its life.

NBV= 400-80= 320

Gain from sale=Cash proceeds(X)- cost- accumulated depreciation

50= X- 400-80

50= X- 320

X= 370